solve each equation for solutions over the interval [0 degrees, 360 degrees)
41. tan2 0 + 4 tan 0 + 2 = 0
Answer #1

Page 200 tam o + 4 tance t 2 foy20) = 0 het tand-u u²+uu+2=0 u= -bt db² yac - - 4 d 16-4(2)(6) -- yt do 20 u= -4 = 282 - 2 -